OpenStudy (anonymous):

differentiate f(x) = [ ln x ]^ 3 i applied the chain rule to obtain 3lnx^2...... do i have to do something else because of absolute value?

OpenStudy (anonymous):

chain rule in this case would be aslo require the derivative of what's in the brackets: \[f(x) = (lnx)^3\] \[f'(x) = 3(lnx)^2*(\frac{1}{x})\]

OpenStudy (anonymous):

ahh i see, the ln is one over x.....thank you !

OpenStudy (anonymous):

yeah, so you bring the exponent out front, keep what's in the brackets, reduce the exponent, and then multiply by the derivative of what's in the brackets

OpenStudy (anonymous):

would there ever be a scenario like [lnx^2]^3 and would the derivative be 6(lnx)^2*(1/x^2)

OpenStudy (anonymous):

ohcorrection *(1/2x)

OpenStudy (anonymous):

yes, that is the chain rule with in a chain rule

OpenStudy (anonymous):

but the derivative of lnx is still be 1/x...i have to go but can dscribe it better tomorrow if your around

OpenStudy (anonymous):

yes i probably will be..... thank you for your help